# Mamba-in-Mamba: Centralized Mamba-Cross-Scan in Tokenized Mamba Model for Hyperspectral Image Classification

### v1

@@ Abstract
Hyperspectral image (HSI) classification is pivotal in the remote sensing (RS) field, particularly with the advancement of deep learning techniques. Sequential models, adapted from the natural language processing (NLP) field—such as Recurrent Neural Networks (RNNs) and Transformers—have been tailored to this task, offering a unique viewpoint. However, several challenges persist: 1) RNNs struggle with centric feature aggregation and are sensitive to interfering pixels, 2) Transformers require significant computational resources and often underperform with limited HSI training samples, and 3) Current scanning methods for converting images into sequence-data are simplistic and inefficient. In response, this study introduces the innovative Mamba-in-Mamba (MiM) architecture for HSI classification, the first attempt of deploying State Space Model (SSM) in this task. The MiM model includes: 1) A novel centralized Mamba-Cross-Scan (MCS) mechanism for transforming images into sequence-data, 2) A Tokenized Mamba (T-Mamba) encoder that incorporates a Gaussian Decay Mask (GDM), a Semantic Token Learner (STL), and a Semantic Token Fuser (STF) for enhanced feature generation and concentration, and 3) A Weighted MCS Fusion (WMF) module coupled with a Multi-Scale Loss Design to improve decoding efficiency. Experimental results from three public HSI datasets with fixed and disjoint training-testing samples demonstrate that our method outperforms existing baselines and state-of-the-art approaches, highlighting its efficacy and potential in HSI applications. The code will be available at https://github.com/zhouweilian1904/Mamba-in-Mamba.
@@ I Introduction
Hyperspectral imaging (HSI) is a sophisticated remote sensing (RS) technique that captures images across numerous dense and continuous spectral bands. This method spans the electromagnetic spectrum from visible to infrared wavelengths, often utilizing satellites or unmanned aerial vehicles (UAVs) [1]. Unlike traditional imaging methods, HSI provides detailed spectral information for each pixel, enabling a more precise characterization of materials and objects on the Earth’s surface [2]. The comprehensive data obtained from HSI makes it invaluable in various applications including agriculture [3], environmental monitoring [4], geology [5], urban planning [6], and defense and security [7]. Given the broad utility of this technology, accurate HSI classification, which assigns a label to each pixel based on its spectral signature, is critical. This classification task is not only essential for enhancing the effectiveness of these applications but has also become a focus of research within the RS community [8].
Figure 1: Conceptual diagram of the Mamba-in-Mamba (MiM) model. Initially, four distinct types of centralized Mamba-Cross-Scan (MCS) are employed to convert a cropped HSI patch into sequential data. It is then fed into the Tokenized Mamba (T-Mamba) encoders designed to process sequences effectively. To adjust the impact of each MCS type, a Weighted MCS Fusion (WMF) is conducted, ensuring a balanced integration of the features extracted from each MCS type. Finally, the outputs from cascaded down-sampling-driven MiM layers are fed into a decoder for classification.
@@
Recently, structured state space sequence (S4) models [39] and the Mamba architecture [40, 41, 42, 43] are brought into the spotlight. These models are adept at handling long-sequence data and can scale linearly with sequence length. The Mamba architecture, in particular, builds on S4 by adding a selective-scan mechanism that adaptively attentions on relevant inputs, showing superior computational efficiency compared to Transformers. Although the Mamba-based methods has shown promising results in some visual tasks [44, 45, 47, 46, 48], its potential for HSI classification has yet to be explored. Therefore, there is a critical need to develop a suitable Mamba model specifically designed for HSI classification.
To this end, in this study, we introduce a innovative approach—the Mamba-in-Mamba (MiM) architecture—specifically designed for HSI classification. This method marks the first application of the visual Mamba framework in this task. The MiM model comprises several innovative components: 1) Centralized Mamba-Cross-Scan (MCS): This new scanning mechanism transforms images into diverse half-directional sequences, offering a more efficient, lightweight, and rational approach, 2) Tokenized Mamba (T-Mamba) Encoder: Integrates a Gaussian Decay Mask (GDM), a Semantic Token Learner (STL), and a Semantic Token Fuser (STF) to significantly enhance representative feature generation and concentration, and 3) Weighted MCS Fusion (WMF) module: Works in conjunction with a Multi-Scale Loss Design to boost decoding efficiency and overall model performance. The conceptual framework of this study is illustrated in Fig. 1.
@@
    For the first time, we explore the potential applications of the vision Mamba architecture for HSI classification task thereby achieving the better performances and efficiency.
@@
    To fully explore the ability of Mamba, we propose a novel T-Mamba encoder with GDM, STL, and STF to make it more suitable for the HSI classification.
@@
    The proposal shows very competitive and even SOTA performance on three datasets with fixed and disjoint training-testing samples, demonstrating the feasibility.
@@ II-A Structured State Space Sequence Model (S4)
Inspired by continuous system theory [49], SSMs are commonly considered linear time-invariant systems that map a signal $\mathbf{x}(t)\in\mathbb{R}^{\mathcal{D}}$ to $\mathbf{y}(t)\in\mathbb{R}^{\mathcal{D}}$ through a latent state $\mathbf{h}(t)\in\mathbb{R}^{\mathcal{N}}$. Here, $t$ represents the time step, $\mathcal{D}$ is the initial dimension, and $\mathcal{N}$ is the dimension of the state.
@@
where $\mathbf{A}\in\mathbb{R}^{\mathcal{N}\times\mathcal{N}}$ is the evolution parameter, $\mathbf{B}\in\mathbb{R}^{\mathcal{N}\times\mathcal{D}}$ correctly aligns dimensions for input-to-state transformation, and $\mathbf{C}\in\mathbb{R}^{\mathcal{D}\times\mathcal{N}}$ is the projection parameter. $\mathbf{D}\in\mathbb{R}^{1}$, which typically supports skip connections, ensures that $\mathbf{y}(t)\in\mathbb{R}^{\mathcal{D}}$.
S4 is a discrete version of a continuous system, developed to address the significant challenges posed when integrating continuous systems into deep learning algorithms. A timescale parameter $\mathbf{\Delta}\in\mathbb{R}^{\mathcal{D}}$ is used to transform the continuous parameters $\mathbf{A}$ and $\mathbf{B}$ into discrete parameters $\mathbf{\overline{A}}\in\mathbb{R}^{\mathcal{N}\times\mathcal{N}}$ and $\mathbf{\overline{B}}\in\mathbb{R}^{\mathcal{N}\times\mathcal{D}}$. The transformation commonly employs a zero-order hold (ZOH) method [^1] https://en.wikipedia.org/wiki/Zero-order_hold, defined as:
@@
Figure 2: The graphic illustration of S4 model.
@@ II-B Selective Scan-Based S4 Model (S6)
1:  Input: Input HSI data $\mathbf{X}\in\mathbb{R}^{\mathcal{H}\times\mathcal{W}\times\mathcal{C}}$, ground truth $\mathbf{L}\in\mathbb{R}^{\mathcal{H}\times\mathcal{W}}$; patch size $p$; fixed training samples.
@@
5:  Apply principle component analysis (PCA) on $\mathbf{X}\in\mathbb{R}^{\mathcal{H}\times\mathcal{W}\times\mathcal{C}}$ for dimension reduction, getting $\mathcal{C}=60$.
@@
7:  Perform initial pixel embedding on $\textbf{X}^{(i,\,j)}$ from feature dimension $\mathcal{C}$ to $\mathcal{C}_{2}$.
8:  Apply four types of MCS on $\textbf{X}^{(i,\,j)}$, getting sequences, $\mathbf{S}_{1}^{(i,\,j)}$, $\mathbf{S}_{2}^{(i,\,j)}$, $\mathbf{S}_{3}^{(i,\,j)}$, and $\mathbf{S}_{4}^{(i,\,j)}$ (Eqs. 13a-18a).
9:  MiM model:
10:  for $m=1$ to $4$ in $\mathbf{S}_{m}^{(i,\,j)}$ do
11:     Split $\mathbf{S}_{m}^{(i,\,j)}$ into sub-sequences, $\mathbf{s}_{f}$ and $\mathbf{s}_{b}$ (Eq. 20).
12:     T-Mamba encoder:
13:     for Both forward pass $\mathbf{s}_{f}$ and backward pass $\mathbf{s}_{b}$ do
14:        Feed sub-sequences $\mathbf{s}_{f}$, $\mathbf{s}_{b}$ into respective Conv1d and S6 model, obtaining $\widetilde{\mathbf{s}_{f}}$ and $\widetilde{\mathbf{s}_{b}}$ (Eqs. 21a-22b).
15:        Apply GDMs to $\widetilde{\mathbf{s}_{f}}$ and $\widetilde{\mathbf{s}_{b}}$, resulting in $\widehat{\mathbf{s}_{f}}$ and $\widehat{\mathbf{s}_{b}}$ (Eqs. 23a and 23b).
16:        Merge two features $\widehat{\mathbf{s}_{f}}$ and $\widehat{\mathbf{s}_{b}}$, getting $\widehat{\mathbf{s}}$ (Eq. 24).
17:        Deploy STL on $\widehat{\mathbf{s}}$ to generate down-sampled semantic tokens $\mathbf{u}$ (Eq. 25).
18:        Apply STF on $\mathbf{u}$ and original feature $\mathbf{z}$ to generate enhanced tokens $\mathbf{\widehat{u}}$ (Eq. 26).
19:        Apply a residual connection between $\mathbf{\widehat{u}}$ and pooled $\mathbf{S}_{m}^{(i,\,j)}$, producing current output $\mathbf{o}_{m}^{(i,\,j)}$ (Eq. 27).
20:        Activate $\mathbf{o}_{m}^{(i,\,j)}$ using the Tanh function.
21:     end for
@@
23:  Merge four outputs, $\mathbf{o}_{1}^{(i,\,j)}$, $\mathbf{o}_{2}^{(i,\,j)}$, $\mathbf{o}_{3}^{(i,\,j)}$, and $\mathbf{o}_{4}^{(i,\,j)}$, by WMF for weighted fusion, recorded as current scale ($p$) feature $\mathbf{O}_{p}^{(i,\,j)}$ (Eq. 48).
24:  Iterate MiM models on $\mathbf{o}_{1}^{(i,\,j)}$, $\mathbf{o}_{2}^{(i,\,j)}$, $\mathbf{o}_{3}^{(i,\,j)}$, and $\mathbf{o}_{4}^{(i,\,j)}$, until the patch size $p$ is reduced to $1$ (Eq. 51).
25:  Train the model using a decoder and a multi-scale loss function (Eq. 52).
@@ II-C Mamba-Based Works
As a cutting-edge foundational model, Mamba has recently gained popularity. There have been various implementations exploring its applicability to diverse tasks, akin to the encoders used in RNNs and Transformers.
Zhu et al. [45] introduced the first Vision Mamba (ViM) architecture for general vision tasks such as object detection and semantic segmentation. ViM utilizes bidirectional scans for image patch positioning before processing with the Mamba encoder. Following this, Chen et al. [46] applied the ViM strategy to enhance Mamba with a dynamic multi-path activation mechanism, achieving superior performance across various RS image classification datasets. Liu et al. [47] developed the Visual SSM (VMamba), which employs cross-scans to convert non-causal visual images into ordered patch sequences for tasks like image classification and object detection. Additionally, Chen et al. [48] extended VMamba’s applications to RS change detection tasks, which include binary change detection, semantic change detection, and building damage assessment. In their approach, VMamba serves as the sequential encoder, modeling spatio-temporal relationships. Concurrently, Zhao et al. [53] followed vanilla Mamba and proposed RS-Mamba for dense prediction in large RS images, introducing the Omni-directional Selective Scan Module (OSSM) for multi-directional feature aggregation.
However, to our knowledge, no studies have yet developed vision Mamba for HSI classification. The suitability of the vanilla Mamba model for HSI classification, its effectiveness compared to RNNs and Transformers, and the potential design of a Mamba architecture optimized for HSI remain open questions. Therefore, we aim to explore these aspects and propose a tailored HSI-driven Mamba architecture.
Figure 3: The overall structure. It begins with a cropped HSI patch that undergoes a pixel embedding stage. It is then divided into four types of sequences by MCS, as shown in Fig. 5. Each scan type is dependently processed by a specially designed T-Mamba encoder, as detailed in Fig. 6, to extract contextual features. The outputs from the four T-Mamba encoders are combined through a WMF module with dynamic weighted scan fusion. Concurrently, fused feature at current scale is recorded for final multi-scale decoding. This process is repeated, with the feature sizes being progressively down-sampled until they reach a spatial size of $1$. Finally, the accumulated multi-scale features from each iteration are compiled to feed into a decoder for model training.

### v3

@@ Abstract
Hyperspectral image (HSI) classification is pivotal in the remote sensing (RS) field, particularly with the advancement of deep learning techniques. Sequential models, adapted from the natural language processing (NLP) field—such as Recurrent Neural Networks (RNNs) and Transformers—have been tailored to this task, offering unique viewpoints. However, they face several problems: 1) RNNs struggle with aggregating central features and are sensitive to interfering pixels; 2) Transformers require extensive computational resources and often underperform with limited HSI training samples. To address the issues associated with RNNs and Transformers, recent advances have introduced State Space Models (SSM) and Mamba, known for their lightweight and parallel scanning capabilities within linear sequence processing, providing a balance between the two. However, the application of existing Mamba-based models to HSI classification has not been fully explored, and the suitability of these models for this task remains an open question. Furthermore, the current methods for converting images into sequence data in Mamba-based models are simplistic and inefficient. In response, this study introduces the innovative Mamba-in-Mamba (MiM) architecture for HSI classification, marking the pioneering deployment of Mamba in this field. The MiM model includes: 1) A novel centralized Mamba-Cross-Scan (MCS) mechanism for transforming images into sequence data; 2) A Tokenized Mamba (T-Mamba) encoder that incorporates a Gaussian Decay Mask (GDM), a Semantic Token Learner (STL), and a Semantic Token Fuser (STF) for enhanced feature generation and concentration; and 3) A Weighted MCS Fusion (WMF) module coupled with a Multi-Scale Loss Design to improve decoding efficiency. Experimental results from four public HSI datasets with fixed and disjoint training-testing samples demonstrate that our method outperforms existing baselines and state-of-the-art approaches, highlighting its efficacy and potential in HSI applications.
@@ I Introduction
Hyperspectral imaging (HSI) is a sophisticated remote sensing (RS) technique that captures images across numerous dense and continuous spectral bands. This method spans the electromagnetic spectrum from visible to infrared wavelengths, often utilizing satellites or unmanned aerial vehicles (UAVs) [1]. Unlike traditional imaging methods, HSI provides detailed spectral information for each pixel, enabling a more precise characterization of materials [2] and objects on the Earth’s surface [3]. The comprehensive data obtained from HSI makes it invaluable in various applications including agriculture [4], environmental monitoring [5], geology [6], urban planning [7], and defense and security [8]. Given the broad utility of this technology, accurate HSI classification, which assigns a label to each pixel based on its spectral signature, is critical. This classification task is not only essential for enhancing the effectiveness of these applications but has also become a focus of research within the RS community [9].
Figure 1: Conceptual diagram of the Mamba-in-Mamba (MiM) model. Initially, four distinct types of centralized Mamba-Cross-Scan (MCS) are employed to convert a cropped HSI patch into sequential data. It is then fed into four shared Tokenized Mamba (T-Mamba) encoders designed to process sequences effectively. To adjust the impact of each MCS type, a Weighted MCS Fusion (WMF) is conducted, ensuring a balanced integration of the features extracted from each MCS type. Finally, the outputs from cascaded down-sampling-driven MiM layers are fed into a decoder for classification.
@@
Recently, structured state space sequence (S4) models [46] and the Mamba architecture [47, 48, 49, 50] are brought into the spotlight. These models are adept at handling long-sequence data and can scale linearly with sequence length. The Mamba architecture, in particular, builds on S4 by adding a selective-scan mechanism that adaptively attentions on relevant inputs, showing superior computational efficiency compared to Transformers. Although the Mamba-based methods has shown promising results in some visual tasks [51, 52, 54, 53, 55, 56] (e.g. Vision Mamba (ViM), Visual Mamba (VMamba), etc), its potential for HSI classification has yet to be fully explored. Therefore, there is a critical need to develop a suitable vision-based Mamba model specifically designed for HSI classification.
To this end, in this study, we introduce a innovative approach—the Mamba-in-Mamba (MiM) architecture—specifically designed for HSI classification. This method marks the pioneering application of the ViM in this task. The MiM model comprises several innovative components: 1) Centralized Mamba-Cross-Scan (MCS): This new scanning mechanism transforms images into diverse half-directional sequences, offering a more efficient, lightweight, and rational approach, 2) Tokenized Mamba (T-Mamba) Encoder: Integrates a Gaussian Decay Mask (GDM), a Semantic Token Learner (STL), and a Semantic Token Fuser (STF) to significantly enhance representative feature generation and concentration, and 3) Weighted MCS Fusion (WMF) module: Works in conjunction with a Multi-Scale Loss Design to boost decoding efficiency and overall model performance. The conceptual framework of this study is illustrated in Fig. 1.
@@
    We explore the potential applications of the improved ViM architecture for HSI classification task thereby achieving the better performances and efficiency.
@@
    To fully explore the ability of ViM, we propose a novel T-Mamba encoder with GDM, STL, and STF to make it more suitable for the HSI classification.
@@
    The proposal shows very competitive and even SOTA performance on four datasets with fixed and disjoint training-testing samples, demonstrating the feasibility.
@@ II-A Structured State Space Sequence Model (S4)
Inspired by continuous system theory [57], SSMs are commonly considered linear time-invariant systems that map a signal $\mathbf{x}(t)\in\mathbb{R}^{1\times\mathcal{D}}$ to $\mathbf{y}(t)\in\mathbb{R}^{1\times\mathcal{D}}$ through a latent state $\mathbf{h}(t)\in\mathbb{R}^{1\times\mathcal{N}}$. Here, $t$ represents the time step, $\mathcal{D}$ is the initial dimension, and $\mathcal{N}$ is the dimension of the state.
@@
where $\mathbf{A}\in\mathbb{R}^{\mathcal{N}\times\mathcal{N}}$ is the evolution parameter, $\mathbf{B}\in\mathbb{R}^{\mathcal{N}\times\mathcal{D}}$ correctly aligns dimensions for input-to-state transformation, and $\mathbf{C}\in\mathbb{R}^{\mathcal{D}\times\mathcal{N}}$ is the projection parameter. $\mathbf{D}\in\mathbb{R}^{1\times\mathcal{D}}$, which typically supports skip connections, ensures that $\mathbf{y}(t)\in\mathbb{R}^{1\times\mathcal{D}}$.
S4 is a discrete version of a continuous system, developed to address the significant challenges posed when integrating continuous systems into deep learning algorithms. A timescale parameter $\mathbf{\Delta}\in\mathbb{R}^{1\times\mathcal{D}}$ is used to transform the continuous parameters $\mathbf{A}$ and $\mathbf{B}$ into discrete parameters $\mathbf{\overline{A}}\in\mathbb{R}^{\mathcal{N}\times\mathcal{N}}$ and $\mathbf{\overline{B}}\in\mathbb{R}^{\mathcal{N}\times\mathcal{D}}$. The transformation commonly employs a zero-order hold (ZOH) method [^1] https://en.wikipedia.org/wiki/Zero-order_hold, defined as:
@@
Figure 2: The graphic illustration of S4 model. $\mathbf{x}(t)$ is a input feature (i.e. scalar or vector) at time step $t$. $\mathbf{h}(t)$ represents the latent state feature. $\mathbf{\overline{A}}$, $\mathbf{\overline{B}}$, $\mathbf{C}$, and $\mathbf{D}$ represents the transformation matrices.
@@ II-B Selective Scan-Based S4 Model (S6)
1:  Input: Input HSI data $\mathbf{X}\in\mathbb{R}^{\mathcal{H}\times\mathcal{W}\times\mathcal{C}_{0}}$, ground truth $\mathbf{L}\in\mathbb{R}^{\mathcal{H}\times\mathcal{W}}$; patch size $p$; fixed training samples.
@@
5:  Apply principal component analysis (PCA) on $\mathbf{X}\in\mathbb{R}^{\mathcal{H}\times\mathcal{W}\times\mathcal{C}}$ for dimension reduction, getting $\mathcal{C}_{1}$.
@@
7:  Perform depth-wise convolution on $\textbf{X}^{(i,\,j)}$ from feature dimension $\mathcal{C}_{1}$ to $\mathcal{C}_{2}$, generating local features.
8:  Perform point-wise convolution on $\textbf{X}^{(i,\,j)}$ then, serving as dynamic positional embedding.
9:  Apply four types of MCS on $\textbf{X}^{(i,\,j)}$, getting sequences, $\mathbf{S}_{1}^{(i,\,j)}$, $\mathbf{S}_{2}^{(i,\,j)}$, $\mathbf{S}_{3}^{(i,\,j)}$, and $\mathbf{S}_{4}^{(i,\,j)}$ (Eqs. 14a-19a).
10:  MiM model:
11:  for $m=1$ to $4$ in $\mathbf{S}_{m}^{(i,\,j)}$ do
12:     Split $\mathbf{S}_{m}^{(i,\,j)}$ into sub-sequences, $\mathbf{s}_{f}$ and $\mathbf{s}_{b}$ (Eq. 21).
13:     T-Mamba encoder:
14:     for Both forward pass $\mathbf{s}_{f}$ and backward pass $\mathbf{s}_{b}$ do
15:        Feed sub-sequences $\mathbf{s}_{f}$, $\mathbf{s}_{b}$ into respective Conv1d and S6 model, obtaining $\widetilde{\mathbf{s}_{f}}$ and $\widetilde{\mathbf{s}_{b}}$ (Eqs. 22a-23b).
16:        Apply GDMs to $\widetilde{\mathbf{s}_{f}}$ and $\widetilde{\mathbf{s}_{b}}$, resulting in $\widehat{\mathbf{s}_{f}}$ and $\widehat{\mathbf{s}_{b}}$ (Eqs. 24a and 24b).
17:        Merge two features $\widehat{\mathbf{s}_{f}}$ and $\widehat{\mathbf{s}_{b}}$, getting $\widehat{\mathbf{s}}$ (Eq. 25).
18:        Deploy STL on $\widehat{\mathbf{s}}$ to generate down-sampled semantic tokens $\mathbf{u}$ (Eq. 26).
19:        Apply STF on $\mathbf{u}$ and original feature $\mathbf{z}$ to generate enhanced tokens $\mathbf{\widehat{u}}$ (Eq. 27).
20:        Apply a residual connection between $\mathbf{\widehat{u}}$ and pooled $\mathbf{S}_{m}^{(i,\,j)}$, producing current output $\mathbf{o}_{m}^{(i,\,j)}$ (Eq. 28).
21:        Activate $\mathbf{o}_{m}^{(i,\,j)}$ using the Tanh function.
@@
23:  end for
24:  Merge four outputs, $\mathbf{o}_{1}^{(i,\,j)}$, $\mathbf{o}_{2}^{(i,\,j)}$, $\mathbf{o}_{3}^{(i,\,j)}$, and $\mathbf{o}_{4}^{(i,\,j)}$, by WMF for weighted fusion, recorded as current scale ($p$) feature $\mathbf{O}_{p}^{(i,\,j)}$ (Eq. 49).
25:  Iterate MiM models on $\mathbf{o}_{1}^{(i,\,j)}$, $\mathbf{o}_{2}^{(i,\,j)}$, $\mathbf{o}_{3}^{(i,\,j)}$, and $\mathbf{o}_{4}^{(i,\,j)}$, until the patch size $p$ is reduced to $1$ (Eq. 52).
26:  Train the model using a decoder (MLP) and a multi-scale loss function (Eq. 53).
@@ II-C Mamba-Based Works
As a cutting-edge foundational model, Mamba has recently gained popularity. There have been various implementations exploring its applicability to diverse tasks.
Zhu et al. [52] pioneered the ViM architecture for general vision tasks such as object detection and semantic segmentation, employing bidirectional scans for precise image patch positioning prior to processing with the Mamba encoder. Building on this framework, Chen et al. [53] enhanced the Mamba with a dynamic multi-path activation mechanism, significantly improving performance across various RS image classification datasets. Concurrently, Liu et al. [54] developed VMamba, which utilizes 2D-Selective-Scan (SS2D) to transform non-causal visual images into ordered patch sequences for applications in image classification and object detection. Further extending this methodology, Chen et al. [55] applied VMamba to RS change detection tasks, encompassing binary change detection, semantic change detection, and building damage assessment, with VMamba serving as the sequential encoder to model spatio-temporal relationships. Additionally, Zhao et al. [61] introduced RS-Mamba for dense prediction in extensive RS images, incorporating the Omni-directional Selective Scan Module (OSSM) for feature aggregation. In the HSI classification domain, Jia et al. [62] adapted the standard SSM-based model, SpectralMamba, for spectral domain processing, while Yang et al. [63] combined Mamba’s spectral processing with CNN-based spatial feature extraction for HSI classification. Moreover, Sheng et al. [64] proposed a lightweight spectral-spatial Mamba-convolution model for HSI classification, utilizing spatial unidirectional and spectral bidirectional scanning for enhanced feature alignment.
Despite these advancements, to our knowledge, the application of ViM with multi-directional scanning for HSI classification has not yet been fully developed or explored. Identifying an optimized ViM architecture for HSI remains an open question. Thus, we aim to explore this gap and propose a tailored HSI-driven ViM architecture.
Figure 3: The overall structure. It begins with a cropped HSI patch after PCA and then goes through a depth-wise convolution and a point-wise convolution for local feature generation and dynamic positional embedding. It is then divided into four types of sequences by MCS, as shown in Fig. 5. Each scan type is independently processed by a specially designed T-Mamba encoder without sharing parameters, as detailed in Fig. 6, to extract contextual features. The outputs from the four T-Mamba encoders are combined through a WMF module with dynamic weighted scan fusion. Concurrently, fused feature at current scale is recorded for final multi-scale decoding. This process is repeated, with the feature sizes being progressively down-sampled until they reach a spatial size of $1$. Finally, the accumulated multi-scale features from each iteration are compiled to feed into a decoder for model training.